• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Vignote

    Log entry time 15:58:34 on March17,2007

    Entry number 196315

    keyword=day shift summary


    Day Shift Summary


    08:00-11:06 kin05, Pb#3, ~40uA, runs 1630-1637 , Total C: 1.214 C
    11:15 MCC phones: No stable current for about 25 minutes
    11:17 stop run 1638, no beam. Total C: 1.224 C
    11:40 start run 1639
    11:40-12:35 kin05, Pb#3, ~40uA, runs 1639,1640 , Total C: 1.332 C
    12:35 Ask MCC to go to 20uA
    12:36-12:46 kin05, Pb#3, ~20uA, run 1641, Total: 1.343 C
    12:46 Ask MCC to go to 10uA
    12:47-12:57 kin05, Pb#3, ~10uA, run 1642, Total: 1.348 C
    13:00 Target to empty
    13:02-13:26 Guido performs a BPM pedestal run, run 1644 (1643 run is junk)
    13:30-13.45 R-spectrometer moved to 48.47 degrees (kin03)
    13:45 Ask MCC to perform a harp scan
    13:50 harp scan result is OK. sigma_y = 0.08 +/- 0.03 mm
    13:55 Target to Carbon
    14:00 Ask MCC current to 40uA and raster on.
    14:02 energy spread too high (8.5E-05). Phone to MCC and page Mark Jones.
    14:13 MCC solves problem with the energy spread. Now it is OK.
    14:13-14:23 kin03, C#2, ~40uA, run 1645 (too high prescale factors)
    14:24-15:17 kin03, C#2, ~40uA, runs 1646,1647 , Total C = 0.132 C
    15:20 Target to Pb
    15:23 Ask MCC to go to 10uA
    15:25-15:35 kin03, Pb#3, ~10uA, run 1648 , Total C = 0.006 C
    15:36 Ask MCC to go to 20uA
    15:37-15:47 kin03, Pb#3, ~20uA, run 1649, Total C = 0.018 C
    15:48 Ask MCC to go to 40uA
    15:49-next shift kin03, Pb#3, ~40uA, run 1650